media%2F835%2F83558103-dea5-4cac-a2ad-74
Answer #1

(a)

Mass% of NaCl in solution = (Mass of NaCl/Mass of solution)*100 = (4.78/10.25)*100 = 46.63%

(b)

Mass-volume% of NaCl in solution = (Mass of NaCl/Volume of solution)*100 = (4.78/12.34)*100 = 38.74%

(c)

Moles of NaCl = Mass/MW = 4.78/58.44 = 0.0818

Molarity = Moles/Volume in liters = 0.0818/0.01234 = 6.63 M